If a * b = (a^2 - b^2) / (a – b), determine the value of
(a * b) * c where a = 3, b = 4, and c = 5.

Answer by drk(1908)

Here is the original problem:
a * b = (a^2 - b^2) / (a – b)
determine the value of (a * b) * c where a = 3, b = 4, and c = 5.
First,
(a*b) = (3*4) = (3^2-4^2) = 9 - 16 = -7/(3-4) = 7
(7*5) = (7^2 - 5^2) / (7-5) = (49-25)/2 = 24/2 = 12